In a sample of 200, we observe 140 successes. Is this sufficient evidence at the 1% significance level to indicate that the population proportion of successes is at least 65%?

Answer by ewatrrr(24785)


Hi

Ho:  p ≥ .65 claim
Ha:  p < .65
z(140/200) = ( .7 - .65)/sqrt[.65*.35/200] = 1.4825
p-value = P(z > 1.4825) = normalcdf(1.4825,100) = 0.069
Conclusion:: Since the p-value is greater than 1%, fail to reject Ho.
there is sufficient evidence at the 1% significance level to indicate 
that the population proportion of successes is at least 65%

Wish You the Best in your Studies.